OpenStudy (henrietepurina):

Choose the point-slope form of the equation below that represents the line that passes through the point (6, -3) and has a slope of 1/2.

OpenStudy (henrietepurina):

y - 6 = one half(x + 3) y = one halfx - 6 y + 3 = one half(x - 6) x - 2y = 12

OpenStudy (henrietepurina):

help for a medal

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ 6}}\quad ,&{\color{blue}{ -3}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{1}{2} \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (henrietepurina):

i dont understand anything

OpenStudy (jdoe0001):

you're given a point, an ordered pair and the slope, for the equation so you use the point-slope form and plug those in :)

OpenStudy (henrietepurina):

ok thanks

OpenStudy (jdoe0001):

yw

OpenStudy (henrietepurina):

y - 6 = 1/2(x - x^1) ok then what

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ 6}}\quad ,&{\color{blue}{ -3}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{1}{2} \\ \quad \\ y-{\color{blue}{ 6}}={\color{green}{ \cfrac{1}{2}}}(x-{\color{red}{ (-3)}})\qquad \textit{plug in the values}\\ \qquad \uparrow\\ \textit{point-slope form} \\ \quad \\ y-6=\cfrac{1}{2}(x+3)\)

OpenStudy (henrietepurina):

ok thanks a lot, heres a medal

OpenStudy (jdoe0001):

yw